OpenStudy (anonymous):

Find the vertical asymptotes, if any, and the horizontal or oblique asymptote, if any, of y= (x^2 -9)/(x^2 -4) The vertical asymptote are x=_____ OR There is no vertical asymptotes The horizontal asymptotes are y=______ OR There are no horizontal asymptotes

OpenStudy (cwrw238):

vertical asymptotes are values of x where the denominator = zero there are 2 vertical asymptotes

OpenStudy (cwrw238):

as function ---> 1 x -->> +/- infinity so what is horizontal asymptote?

OpenStudy (cwrw238):

this should help

OpenStudy (anonymous):

Confuse...

OpenStudy (campbell_st):

can you factor the denominator..?

OpenStudy (cwrw238):

vertical asymptote: x^2 - 4 = 0 - solve this bay factorising to give 2 values of x

OpenStudy (anonymous):

okay so it will be (x-3)(x+3)/(x-2)(x+2)=y

OpenStudy (campbell_st):

the simply way to find the oblique or horizontal asymptote ( in this case) is to simplify \[\frac{x^2}{x^2} = ?\] normally it takes a little polynomial division.... but can be found by looking at the leading terms...

OpenStudy (campbell_st):

thats correct.. so the denominator can't be zero so the vertical asymptotes occur when x - 2 = 0 and x + 2 = 0 solve for x to find the asymptotes

OpenStudy (anonymous):

it will be 2 and -2

OpenStudy (campbell_st):

thats correct... now you have the vertical asymptotes. go bale and simplify for the horizontal...

OpenStudy (anonymous):

wait how to find for horizontal?

OpenStudy (anonymous):

plug in for x right?

OpenStudy (campbell_st):

you can use polynomial division or in this question because the leading terms have the same power just simplify \[\frac{x^2}{x^2} = ?\] whats the value of ?

OpenStudy (anonymous):

I got 11/6 and 7/2 when t plug in x=-2 and 2 into the original equation.

OpenStudy (anonymous):

would that be right?

OpenStudy (jdoe0001):

when numerator and denominator's leading term have the same degree, the horizontal asymptote would be at their coefficient's division -> n/d

OpenStudy (campbell_st):

don't plug in any values.... you don't need to just whats the simplified value of \[\frac{x^2}{x^2} = 1\] so the oblique or horizontal asymptote is x = 1

OpenStudy (anonymous):

okay so do you use that equation to find for the horizontal. would it always = to 1

OpenStudy (campbell_st):

nope... it requires the numerator to be divided by the denominator... to find the oblique asymptote... its just in this question its nice and neat and equal to 1.

OpenStudy (jdoe0001):

there should be an oblique asymptote, if the numerator degree were 1 higher than the denominator :S, in this case, they're both the same

OpenStudy (anonymous):

confuse so its not 1 for the oblique?

OpenStudy (jdoe0001):

no, it means it has no oblique asymptote

OpenStudy (campbell_st):

here is what you need vertical asymptotes at x = -2, 2 horizontal asymptote at y = 1 thats all you need

OpenStudy (jdoe0001):

so as campbell_st said, 1 is correct, that is y=1, but that's just a straight line, no an oblique one

OpenStudy (campbell_st):

its oblique with a slope of zero....

OpenStudy (anonymous):

okay thank you guys for clearing it for me!